Simplify: 0 x 10 2 Simplify: 0 x 10 2 Simplify: 0 x 10 2 Simplify: 0 x 10 2

Simplify:
0 x 102
Simplify:
0 x 102
Simplify:
0 x 102
Simplify:
0 x 102
 

Grade:12th pass

2 Answers

Harshit Singh
2 years ago
Dear Student

The above question can be written as,
= 0 × 10 × 10
= 0 × 100
=0

Thanks
Shraddha Singh
2 years ago
Dear Student,
 
We know that any number multiplied to 0 gives 0.
Hence ,
0 × 10²
= 0 × 10 × 10 = 0
 
Hope it helps!
Thank you ✌️